A parallel-plate capacitor has plates of area 2.88×10−4 m2 .

Part A.

What plate separation is required if the capacitance is to be 1660 pF ? Assume that the space between the plates is filled with air? (Dielectric constant for air is 1.00059)

Express your answer using three significant figures.d= ? μm

Part B

What plate separation is required if the capacitance is to be 1660 pF ? Assume that the space between the plates is filled with paper. (Dielectric constant for paper is 3.7)

Express your answer using two significant figures. d= ? μm

Please show all of your steps, thank you!
Answer #1

C = epsilon _not * A / d

a)

1660*10^-12 = 8.85*10^-12* 2.88*10^-4 / d

d = 1.535 um

========

b)

1660 *10^-12 = 3.7* 8.85*10^-12* 2.88*10^-4/ d

d = 5.7 um

=======

Comment in case any doubt, will reply for sure.. Goodluck